Hi!


I have a question on a probability problem. You draw a fi ve card hand from a deck of 52 cards and nd that you are 2 cards away from a Flush. What is the probability that, if you remove the two bad cards, you will pull the correct cards from the deck to make a Flush?

Answer by shree840(260)

How many players are there. Are you playing with one deck of cards?
5cards are taken out.
prob of next right card =1/47
prob of second right card=1/46
combined =(1/46)(1/47) ans